# RapidPoseTriangulation: Multi-view Multi-person Whole-body Human Pose Triangulation in a Millisecond

### v4

@@ 1 Introduction
Figure 1: Example of a multi-person whole-body pose estimation from multiple camera views in a volleyball game (from the egohumans dataset [khirodkar2023ego]). On top, the full image of one camera with the projections of all the detected poses, on the bottom-left a zoom-in on one player, and on the bottom-right her detected 3D pose.
@@
Recent advances in deep learning have dramatically improved the accuracy of pose estimation from individual camera views. However, the task of efficiently combining these multiple 2D estimates into accurate 3D poses, particularly for multiple people in real-time scenarios, remains an active area of research. Current approaches often show notable performance drops when applied to setups they were not trained on, and normally are too slow for real-time usage. Furthermore, as applications become more advanced, there is an increasing demand for a more detailed whole-body pose estimation, including facial expressions and especially hand gestures (see Figure 1).
@@ 2 Related Work
Most approaches address the problem of human pose estimation in two separate steps. At first, 2D poses are predicted for each image, and in the second step, these are fused to estimate 3D poses. The basic concepts of the different approaches in the second step can be categorized on whether they use algorithmic strategies or learning-based methods. Some of them also make use of temporal information from previous frames to improve the results.
@@
On the side of the learning-based methods, VoxelPose [voxelpose] was one of the first concepts, building upon the research of Iskakov et al. [iskakov2019learnable] and extending it to multi-person estimations. Its method projects joint heatmaps from 2D images into a 3D voxelized space, estimates a central point for each individual using a convolution network, extracts a focused cube surrounding each person center, and computes the joint positions using a second net. Faster-VoxelPose [fastervoxelpose] enhanced this technique by restructuring the 3D voxel network into several 2D and 1D projections, thereby boosting efficiency. TesseTrack [reddy2021tessetrack] and TEMPO [choudhury2023tempo] incorporated a temporal dimension into the voxel space to refine poses across sequential frames. Alternative approaches such as PRGnet [wu2021graph] employ a graph-based methodology or directly infer 3D poses from 2D features, like in MvP [wang2021mvp]. SelfPose3d [Srivastav_2024_CVPR] adopts the framework of VoxelPose and trains both 2D and 3D networks in a self-supervised manner using randomly augmented 2D poses.
@@
Regarding algorithmic methodologies, mvpose [dong2019fast] solves the estimation problem in two steps. Initially, it identifies corresponding 2D poses across images based on geometric and visual clues, and then triangulates the matching pose groups to calculate the final output. mv3dpose [tanke2019iterative] utilizes a graph-matching approach to assign poses through epipolar geometry and incorporates temporal information to handle missing joint data. PartAwarePose [chu2021part] accelerates the pose-matching process by leveraging poses from preceding frames and applying a joint-based filter to solve keypoint inaccuracies resulting from occlusions. VoxelKeypointFusion [voxkeyfuse] employs a voxel-based triangulation concept to predict 3D joint proposals from overlapping viewpoints. It then uses the joint reprojections to assign them to individuals in the original 2D views, to match them to individual 3D persons before merging them into a final result.
@@
Regarding the generalization of learning-based approaches to new datasets, a direct transfer is typically not evaluated. However, VoxelPose, Faster-VoxelPose, MvP, and TEMPO have implemented a finetuning strategy utilizing synthetic data, though only the first two have also published the source-code for this. The idea is to randomly place 3D poses from another dataset in 3D space, and back-project them onto the camera perspectives of the target setup. After adding some augmentations, the network is then trained to learn the 3D reconstruction process. In the closed-source approach of CloseMoCap [shuai2023reconstructing] this concept was further improved by using a larger 3D pose dataset and additional augmentations. Algorithmic approaches generally evaluated their transferability to a limited number of other datasets, a process that is considerably more straightforward as it doesn’t require any training.
@@ 3.1 Algorithm
Similar to most other approaches, the algorithm can be split into two stages as well, with the first one predicting the 2D poses for each image. For this any 2D pose estimator can be integrated, here RTMPose [jiang2023rtmpose] was used. The second stage can be split into the following steps:
@@
Figure 2: Obtaining 3D proposals. The process starts with 2D detections for each view (a, images from chi3d [fieraru2020three], note the small blue colored false positive). Then, in step (1), all possible pairs between the views are created. In this case this leads to six pairs from the two detections above with the three below. The core joints of all pairs are triangulated into 3D proposals (b, steps (3,4)). Then they are reprojected into the 2D views (c, step (6)), and a distance-based error to the original 2D poses (visualized in black in c) is calculated. As can be seen in the image, the green and pink proposals clearly do not match to their original 2D poses, and get a very high error. The yellow and light-blue poses resulted from the flipped (man with woman) pairs and also have a notable error. All pairs with errors above a threshold are dropped in step (8). Only the remaining red and dark-blue pairs with low enough errors are used for the further steps (9-12).
@@
Optionally, in steps (13-14) the results can be tracked and smoothed over time. A simple distance-based threshold is used to assign poses to existing tracks. Just clipping the joint movements to a maximum speed was preferred to more complex smoothing filters, because it does not introduce response latency to sudden direction changes, which could hurt real-time position accuracy.
@@ 3.2 Method Comparison
There exist several differences to the other algorithmic approaches. In comparison to mvpose [dong2019fast] no person identification by appearance similarities is required, the association is done only by geometric matching. mv3dpose [tanke2019iterative] matches and groups 2D poses by epipolar geometry instead of the 3D triangulation and grouping used here. PartAwarePose [chu2021part] groups the 2D poses by assigning them to the 3D poses from the previous view, and uses a temporal filter to predict future locations from joints using previous observations to filter joints, which on the other hand can lead to prediction errors and costs additional computation time. 4DAssociation [zhang20204d] follows a bottom-up graph-based concept to match body-parts and connections over space and time to assemble the 3D poses afterward. Bridgeman et al [bridgeman2019multi] group the 2D poses to persons by first calculating correspondence costs for all pairs of 2D poses, using the minimal distance between two ray projections, calculate 3D poses and then iteratively calculate costs from the point-to-ray distance to update the 2D-3D matches. The results are further smoothed by a temporal filter. The source-code was not published. QuickPose [zhou2022quickpose] splits the detected 2D skeletons into all plausible part-graphs, calculates a ray-based distance between joints in different views, and uses a clustering algorithm to find matching parts. Chen et al. [Chen_2020_CVPR] iteratively match 2D poses from one view to the already existing 3D persons, and gradually update the matched 3D poses, before continuing with the next view. The last two algorithms are the fastest ones yet, according to the times stated in their papers, but their source-codes were not published, therefore a detailed comparison is not possible. In comparison to VoxelKeypointFusion [voxkeyfuse] no voxel space is used, which results in much faster calculations, and its person association follows a bottom-up instead of the top-down approach used here.
@@ 3.3 Performance
Due to the relatively simple concept, the algorithm is very fast. In a test on the shelf [belagiannis20143d] dataset, which shows up to four persons in five views, the algorithm requires an average triangulation time of only 0.1ms0.1\,ms. The per-step times are listed in Table 1.
@@
While the initial version of the algorithm used OpenCV functions for most calculations, the current version uses standard C++ functions only, because the overhead of the OpenCV functions was higher than the actual calculation time for the relatively small number of points. Instead of OpenCV’s SVD-based triangulation, a simple mid-point triangulation was implemented, which was much faster and almost as accurate in the experiments. Using only the core joints, and repeatedly filtering invalid poses helps to keep the subsequent steps efficient. The runtime is further influenced by the number of joints, which are by default $20$. In the whole-body case that uses $136$ joints, the average runtime increases to 0.4ms0.4\,ms. Depending on the integration of the algorithm, data conversion time can lead to further delays. The algorithm itself is implemented in C++, but an interface to Python (using SWIG) is provided as well.
As shown in Table 2 the proposed solution is multiple times faster than all the other algorithms, which proves its significant efficiency improvements. Note that all times reported in the different papers were measured on different hardware, here a single core of an AMD-7900X was used.
@@
4DAssociation [zhou2022quickpose] 32
PartAwarePose [chu2021part] 10
Bridgeman et al. [bridgeman2019multi] 9.1
Chen et al. [Chen_2020_CVPR] 3.1
QuickPose [zhou2022quickpose] 2.9
@@
Table 2: Time from 2D to 3D poses on shelf [belagiannis20143d] in milliseconds.
@@ 4 Dataset Generalization
The most important part for machine learning models is their generalization to previously unknown setups. Therefore the proposed approach was evaluated on a variety of datasets. The datasets used for the evaluation were human36m [h36m_pami], shelf [belagiannis20143d], campus [belagiannis20143d], mvor [srivastav2018mvor], and panoptic [joo2015panoptic], similar as in other works. Additionally, the datasets chi3d [fieraru2020three], tsinghua [20204DAssociation] and egohumans [khirodkar2023ego] were added as well to allow an even broader evaluation. The metrics used are the same as in VoxelKeypointFusion [voxkeyfuse], and evaluate the same $13$ keypoints (2 shoulders, 2 hips, 2 elbows, 2 wrists, 2 knees, 2 ankles, 1 nose/head) across each dataset. The metrics follow common standards and are described in more detail in [voxkeyfuse]. The FPS were measured on an Nvidia-3090 as well. Due to brevity and readability reasons, for the datasets already evaluated in VoxelKeypointFusion [voxkeyfuse], only the best results/algorithms are compared here.
@@
Table 3: Transfer to human36m [h36m_pami], all other results from [voxkeyfuse].
@@
Table 4: Transfer to shelf [belagiannis20143d], all other results from [voxkeyfuse].
@@
Table 5: Transfer to campus [belagiannis20143d], all other results from [voxkeyfuse].
@@
SimpleDepthPose [simdepose] 74.0    62.0 94.3 113    54.1 96.6 23.5 85.4 37.2
Faster-VoxelPose (synthetic) [voxkeyfuse] 37.9    28.1 45.5 109    29.4 46.1 7.7 61.5 30.0
VoxelKeypointFusion [voxkeyfuse] 54.5    43.9 75.1 128    35.9 76.6 24.2 76.2 11.3
@@
Table 6: Transfer to mvor [srivastav2018mvor] with depth images (first) and without (others).
@@
Table 7: Replication of panoptic [joo2015panoptic] results and transfer without depth, all other results from [voxkeyfuse].
@@ 4.1 Standard datasets
A lot of the MPJPE error in the human36m dataset results from different skeleton definitions. The original labels for example have higher hips, the nose more inside the head, and the ankles more at the heels, than the COCO skeleton definition of the 2D pose estimation network. From a human perspective though, the predictions would often match better to the images than the original labels.
@@
A lot of errors in the panoptic dataset involve the lower body joints, especially the ankles. While most upper body joints have an average error between 1515-25mm25\,$m$, the hips and knees have around 3535-40mm40\,$m$, and the ankles around 50mm50\,$m$. This is mainly caused by the position of the cameras, which often cut off part of the legs. The 2D pose estimation network, however, still predicts those joints at the lower edge of the image (so for example an ankle is predicted at the height of the knee). This leads to a large error in the triangulation step because the 3D joint is then estimated with a wrong depth, which not always can be detected as outlier. For future optimization, this might be better handled if the 2D pose estimation network could predict the visibility of the keypoints as well, so the algorithm could prefer directly visible keypoints over estimated ones.
@@ 4.2 Additional datasets
To allow an even broader comparison, two further datasets were added, following the same scheme as in VoxelKeypointFusion [voxkeyfuse] and using their skelda dataset library.
@@
Table 8: Transfer to chi3d [fieraru2020three]
@@
Table 9: Transfer to tsinghua [20204DAssociation]
@@
The first one is the chi3d [fieraru2020three] dataset, which shows two persons interacting with each other. It has the challenge that the persons are often very close to each other, which makes it hard to distinguish between them. The setup is similar to human36m and also contains four cameras. The results in Table 8 show that generally most persons are detected, with a similar performance as in human36m. Regarding RapidPoseTriangulation, it could be seen that in some cases the 2D pose estimation struggled to correctly detect the joints between very close persons, and the 3D triangulation then did not get enough inputs to correctly triangulate all joints and could miss a person. One could reduce the reprojection matching threshold to better handle such problematic 2D detections, but at the cost of increasing the number of false positives, or alternatively use the tracking information to fill-up missed persons with their last positions. The authors of CloseMoCap [shuai2023reconstructing] also reported very good transfer results on this dataset, reaching a PCK@50 of $94.3$ on average. They also evaluated VoxelPose, Faster-VoxelPose and mvpose on the same dataset and reached much better results for them as well. Besides that, they reported per-joint errors on a subset of the dataset and found that the hip joints had the best results, whereas here the hip joints are among the worst results. In the labels, it though can be seen that the hips are positioned higher on the body than in the COCO skeleton definition. So in conclusion the mismatch is likely caused by parts of the evaluation, or the use of different labels, somewhere in CloseMoCap, but because their source-code was not published, this assumption can not be verified.
The second dataset is tsinghua [20204DAssociation], which shows people moving around in a room, while they are being recorded from six different cameras. To prevent multiple trainings for the synthetic models, only the first of the two sequences was evaluated. The results in Table 9 show that almost all models perform well on this dataset.

### v5

@@ 1 Introduction
Figure 1: Example of a multi-person whole-body pose estimation from multiple camera views in a volleyball game (from the egohumans dataset [16]). On top, the full image of one camera with the projections of all the detected poses, on the bottom-left a zoom-in on one player, and on the bottom-right her detected 3D pose.
@@
Recent advances in deep learning have dramatically improved the accuracy of pose estimation from individual camera views. However, the task of efficiently combining these multiple 2D estimates into accurate 3D poses, particularly for multiple people in real-time scenarios, remains an active area of research. Current approaches often show notable performance drops when applied to setups they were not trained on, and are normally too slow for real-time usage [2]. Furthermore, as applications become more advanced, there is an increasing demand for a more detailed whole-body pose estimation, including facial expressions and especially hand gestures (see Figure 1).
@@ 2 Related Work
Most approaches address the problem of human pose estimation in two separate steps. At first, 2D poses are predicted for each image, and in the second step, these are fused to estimate 3D poses. The basic concepts of the different approaches in the second step can be categorized by whether they use algorithmic strategies or learning-based methods. Some of them also make use of temporal information from previous frames to improve the results.
@@
On the side of the learning-based methods, VoxelPose [24] was one of the first concepts, building upon the voxel-based triangulation research of Iskakov et al. [13] and extending it from single-person to multi-person estimations. Its method projects joint heatmaps from 2D images into a 3D voxelized space, estimates a central point for each individual using a convolution network, extracts a focused cube surrounding each person center, and computes the joint positions using a second net. Faster-VoxelPose [28] enhanced this technique by restructuring the 3D voxel network into several 2D and 1D projections, thereby boosting efficiency. TesseTrack [18] and TEMPO [8] added a temporal dimension into the voxel space to refine poses across sequential frames. Alternative approaches such as PRGnet [26] employ a graph-based methodology or directly infer 3D poses from 2D features, like in MvP [25]. SelfPose3d [20] adopts the framework of VoxelPose and trains both 2D and 3D networks in a self-supervised manner using randomly augmented 2D poses.
@@
Regarding algorithmic methodologies, mvpose [10] solves the estimation problem in two steps. Initially, it identifies corresponding 2D poses across images based on geometric and visual clues, and then triangulates the matching pose groups to calculate the final output. mv3dpose [23] utilizes a graph-matching approach to assign poses through epipolar geometry and incorporates temporal information to handle missing joint data. PartAwarePose [9] accelerates the pose-matching process by leveraging poses from preceding frames and applying a joint-based filter to solve keypoint inaccuracies resulting from occlusions. VoxelKeypointFusion [2] employs a voxel-based triangulation concept to predict 3D joint proposals from overlapping viewpoints. It then uses the joint reprojections to assign them to individuals in the original 2D views, to match them to individual 3D persons before merging them into a final result.
@@
Regarding the generalization of learning-based approaches to new datasets, a direct transfer is typically not evaluated. However, VoxelPose, Faster-VoxelPose, MvP, and TEMPO have implemented a finetuning strategy utilizing synthetic data, though only the first two have also published the source-code for this. The idea is to randomly place 3D poses from another dataset in 3D space, and back-project them onto the camera perspectives of the target setup. After adding some augmentations, the network is then trained to learn the 3D reconstruction process. In the closed-source approach of CloseMoCap [19] this concept was further improved by using a larger 3D pose dataset and additional augmentations. Algorithmic approaches generally evaluated their transferability to a limited number of other datasets, a process that is considerably more straightforward as it doesn’t require any training.
@@ 3.1 Algorithm
Similar to most other approaches, the algorithm can be split into two stages as well, with the first one predicting the 2D poses for each image. For this any 2D pose estimator can be integrated, here RTMPose [14] was used. The second stage can be split into the following steps:
@@
Figure 2: Obtaining 3D proposals. The process starts with 2D detections for each view (a, images from chi3d [11], note the small blue colored false positive in the second image). Then, in step (1), all possible pairs between the views are created. In this case this leads to six pairs from the two detections above with the three below. The core joints of all pairs are triangulated into 3D proposals (b, steps (3,4)). Then they are reprojected into the 2D views (c, step (6)), and a distance-based error to the original 2D poses (visualized in black in c) is calculated. As can be seen in the image, the green and pink proposals clearly do not match to their original 2D poses, and get a very high error. The yellow and light-blue poses resulted from the flipped (man with woman) pairs and also have a notable error. All pairs with errors above a threshold are dropped in step (8). Only the remaining red and dark-blue pairs with low enough errors are used for the further steps (9-12).
@@
Optionally, in steps (13-14) the results can be tracked and smoothed over time. A simple distance-based threshold is used to assign poses to existing tracks. Just clipping the joint movements to a maximum speed was preferred to more complex smoothing filters, because it does not introduce response latency to sudden direction changes, which could hurt real-time-position accuracy.
@@ 3.2 Method Comparison
There exist several differences to the other algorithmic approaches. In comparison to mvpose [10] no person identification by appearance similarities is required, the association is done only by geometric matching. mv3dpose [23] matches and groups 2D poses by epipolar geometry instead of the 3D triangulation and grouping used here. PartAwarePose [9] groups the 2D poses by assigning them to the 3D poses from the previous view, and uses a temporal filter to predict future locations from joints using previous observations to filter joints, which on the other hand can lead to prediction errors and costs additional computation time. 4DAssociation [29] follows a bottom-up graph-based concept to match body-parts and connections over space and time to assemble the 3D poses afterward. Bridgeman et al [4] group the 2D poses to persons by first calculating correspondence costs for all pairs of 2D poses, using the minimal distance between two ray projections, calculate 3D poses and then iteratively calculate costs from the point-to-ray distance to update the 2D-3D matches. The results are further smoothed by a temporal filter. The source-code was not published. QuickPose [30] splits the detected 2D skeletons into all plausible part-graphs, calculates a ray-based distance between joints in different views, and uses a clustering algorithm to find matching parts. Chen et al. [6] iteratively match 2D poses from one view to the already existing 3D persons, and gradually update the matched 3D poses, before continuing with the next view. The last two algorithms are the fastest ones yet, according to the times stated in their papers, but their source-codes were not published, therefore a detailed comparison is not possible. In comparison to VoxelKeypointFusion [2] no voxel space is used, which results in much faster calculations, and its person association follows a bottom-up instead of the top-down approach used here.
@@ 3.3 Performance
Due to the relatively simple concept, the algorithm is very fast. In a test on the shelf [1] dataset, which shows up to four persons in five views, the algorithm requires an average triangulation time of only 0.1ms0.1\,ms. The per-step times are listed in Table 1.
@@
While the initial version of the algorithm used OpenCV functions for most calculations, the current version uses standard C++ functions only, because the overhead of the OpenCV functions was higher than the actual calculation time for the relatively small number of points. Instead of OpenCV’s SVD-based triangulation, a simple mid-point triangulation was implemented, which was much faster and almost as accurate in the experiments. In summary the switch to standard C++ functions led to a significant relative speedup of around 0.1ms0.1\,ms. Using only the core joints, and repeatedly filtering invalid poses helps to keep the subsequent steps efficient. The runtime is further influenced by the number of joints, which are by default $20$. In the whole-body case that uses $136$ joints, the average runtime increases to 0.4ms0.4\,ms. Depending on the integration of the algorithm, data conversion time can lead to further delays. The algorithm itself is implemented in C++, but an interface to Python (using SWIG) is provided as well.
As shown in Table 2 the proposed solution is multiple times faster than all the other algorithms, which proves its significant efficiency improvements. Note that all times reported in the different papers were measured on different hardware (see appendix), here a single core of an AMD-7900X was used, but the main speedup is clearly caused by the more efficient algorithm.
@@
4DAssociation [30] 32
PartAwarePose [9] 10
Bridgeman et al. [4] 9.1
Chen et al. [6] 3.1
QuickPose [30] 2.9
@@
Table 2: Time from 2D to 3D poses on shelf [1] in milliseconds.
@@ 4 Dataset Generalization
The most important part for machine learning models is their generalization to previously unknown setups. Therefore the proposed approach was evaluated on a variety of datasets. The datasets used for the evaluation were human36m [12], shelf [1], campus [1], mvor [21], and panoptic [15], similar as in other works. Additionally, the datasets chi3d [11], tsinghua [29] and egohumans [16] were added as well to allow an even broader evaluation. The metrics used are the same as in VoxelKeypointFusion [2], and evaluate the same $13$ keypoints (2 shoulders, 2 hips, 2 elbows, 2 wrists, 2 knees, 2 ankles, 1 nose/head) across each dataset. The metrics follow common standards and are described in more detail in [2]. The FPS were measured on an Nvidia-3090 as well. Due to brevity and readability reasons, for the datasets already evaluated in VoxelKeypointFusion [2], only the best results/algorithms are compared here.
@@
Table 3: Transfer to human36m [12], all other results from [2].
@@
Table 4: Transfer to shelf [1], all other results from [2].
@@
Table 5: Transfer to campus [1], all other results from [2].
@@
SimpleDepthPose [3] 74.0    62.0 94.3 113    54.1 96.6 23.5 85.4 37.2
Faster-VoxelPose (synthetic) [2] 37.9    28.1 45.5 109    29.4 46.1 7.7 61.5 30.0
VoxelKeypointFusion [2] 54.5    43.9 75.1 128    35.9 76.6 24.2 76.2 11.3
@@
Table 6: Transfer to mvor [21] with depth images (first) and without (others).
@@
Table 7: Replication of panoptic [15] results and transfer without depth, all other results from [2].
@@ 4.1 Standard datasets
Much of the MPJPE error in the human36m dataset results from different skeleton definitions. The original labels for example have higher hips, the nose more inside the head, and the ankles more at the heels, than the COCO [17] skeleton definition of the 2D pose estimation network. From a human perspective though, the predictions would often match better to the images than the original labels.
@@
Many errors in the panoptic dataset involve the lower body joints, especially the ankles. While most upper body joints have an average error between 1515-25mm25\,$m$, the hips and knees have around 3535-40mm40\,$m$, and the ankles around 50mm50\,$m$. This is mainly caused by the position of the cameras, which often cut off part of the legs. The 2D pose estimation network, however, still predicts those joints at the lower edge of the image (so for example an ankle is predicted at the height of the knee). This leads to a large error in the triangulation step because the 3D joint is then estimated with a wrong depth, which cannot always be detected as an outlier. For future optimization, this might be better handled if the 2D pose estimation network could predict the visibility of the keypoints as well, so the algorithm could prefer directly visible keypoints over estimated ones.
@@ 4.2 Additional datasets
To allow an even broader comparison, two further datasets were added, following the same scheme as in VoxelKeypointFusion [2], and using their skelda dataset library.
@@
Table 8: Transfer to chi3d [11]
@@
Table 9: Transfer to tsinghua [29]
@@
The first one is the chi3d [11] dataset, which shows two persons interacting with each other. It has the challenge that the persons are often very close to each other, which makes it hard to distinguish between them. The setup is similar to human36m and also contains four cameras. The results in Table 8 show that generally most persons are detected, with a similar performance as in human36m. Regarding RapidPoseTriangulation, it could be seen that in some cases the 2D pose estimation struggled to correctly detect the joints between very close persons, and the 3D triangulation then did not get enough inputs to correctly triangulate all joints and could miss a person. One could reduce the reprojection matching threshold to better handle such problematic 2D detections, but at the cost of increasing the number of false positives, or alternatively use the tracking information to fill-up missed persons with their last positions. The authors of CloseMoCap [19] also reported very good transfer results on this dataset, reaching a PCK@50 of $94.3$ on average. They also evaluated VoxelPose, Faster-VoxelPose and mvpose on the same dataset and reached much better results for them as well. Additionally, they reported per-joint errors on a subset of the dataset and found that the hip joints had the best results, whereas here the hip joints are among the worst results. In the labels, however, it can be seen that the hips are positioned higher on the body than in the COCO skeleton definition, similar as already in human36m. So in conclusion, the mismatch is likely caused by parts of the evaluation, or the use of different labels, somewhere in CloseMoCap, but because their source-code was not published, this assumption cannot be verified.
The second dataset is tsinghua [29], which shows people moving around in a room, while they are being recorded from six different cameras. To prevent multiple trainings for the synthetic models, only the first of the two sequences was evaluated. The results in Table 9 show that almost all models perform relatively well on this dataset, but still many of them miss a few persons.